1. A bullet is fired at an angle of 30degrees to the horizontal with a velocity of 150m/s calculate the time the bullet takes to attain it's maximum height
1
Expert's answer
2022-03-30T13:49:41-0400

The time required the bullet takes to attain it's maximum height

"t=\frac{v_0\sin\theta}{g}"

"t=\frac{150*\sin30^\circ}{9.8}=7.65\;\rm s"